A germanium crystal with doping of 1018arsenic atoms per cm3

is required. For

initial Ge load of 20 kg, calculate the mass of arsenic to be added. Assume that the

value of k0 remains constant throughout the growth process. Consider the density of

molten germanium to be 5.7 g cm-3

and atomic weight of arsenic 74.934u.


1
Expert's answer
2021-02-24T12:50:59-0500
"V=\frac{m_{Ge}}{\rho_{Ge}}=\frac{20000}{5.7}=3509\ cm^3"

"m_{As}=\frac{N}{N_A}M_{As}=\frac{nV}{N_A}M_{As}\\
m_{As}=\frac{10^{18}\cdot3509}{6.02\cdot10^{23}}74.934=0.437\ g"